You plan to invest $2,000 in a corporate stock fund or in a common fund. Under certain conditions, the yield of the fund is as follows:

Probability Condition Corporate Return Common Fund Return
0.10 Recession -$200 -$150
0.15 Stagnation $100 $80
0.35 Slow Growth $200 $200
0.30 Moderate Growth $400 $500
0.10 Fast Growth $600 $550

Question 8 (1 point)

Compute the expected return for both the corporate and common fund.
Answer #1

expected return of corporate fund =xP(x)=0.1*(-200)+0.15*(100)+0.35*200+0.3*400+0.1*600=$ 245

expected return for common fund =0.1*(-150)+0.15*80+0.35*200+0.3*500+0.1*550=$ 272